Mass of Evaporating Dish & Zinc= 40,7715 g
Mass of Evaporating Dish= 40,2948 g
Mass of Zinc= 0,4767 g
Mass of Evaporating Dish and Zinc Chloride=41,3245 g (1st attempt)
Mass of Zinc Chloride =?????
Mass of Chlorine in Zinc Chloride=?????

HOW TO DETERMINE MASS OF ZINC CHLORIDE AND MASS OF CHLORINE IN ZINC CHLORIDE??? PLEASE HELP
1
Expert's answer
2020-11-14T13:56:31-0500

m(ZnCl2)=41.3245-40.7715=0.553(g)

To find mass of chlorine here the mass percent of Cl should be found.

%(Cl)={2×Ar(Cl)/Mr(ZnCl2)}×100%

Mr(ZnCl2)=65+2×35.5=136(g/mol)

%(Cl)={71/136}×100%=52.2%

m(Cl)=m(ZnCl2)*%(Cl)/100%=0.553×0.522=0.2887(g)